# Instruction-Based Molecular Graph Generation with Unified Text-Graph Diffusion Model

### v1

@@ Introduction
Table 1: Results of the instruction-based molecule generation on ChEBI-20 dataset for both with/without pretraining setting.
Pretraining Method
@@
Recent studies have demonstrated that BERT can be viewed as a one-step [MASK] absorbing diffusion model  (Devlin et al. 2019; Austin et al. 2021), since one-step noising replaces some tokens with [MASK] and denoising performs masked prediction. This shared training objective allows for the effective initialization from pre-trained language models (He et al. 2022). Therefore, to preserve language comprehension, we utilize the pre-trained RoBERTa model (Liu et al. 2019) for initialization.
@@
However, the introduction of novel graph tokens during tokenization renders the direct loading of pre-trained BERT/RoBERTa parameters insufficient for mask prediction on graph tokens. Consequently, our pretraining methods are designed to equip the model with mask prediction capabilities for graph data while retaining its functionality for text data. We collect paired or single-modal data in both text and molecular graph modalities, mask all tokens and edge indices with a 15% probability, and pretrain the model using the masked language prediction method. The effectiveness of this pretraining approach for graph data has been validated by previous studies on graph pretraining (Hu et al. 2020; Hou et al. 2022).
Experiment
Dataset
@@
Table 2: Review of pretraining datasets. Some data listed in BioT5 is unavailable, and we reproduce it with our dataset.
Pretraining Dataset
We collect several open-source datasets for pretraining. For the molecule-text pair data, we select the PubChem-324K dataset (Liu et al. 2023c), which contains approximately 320,000 molecule-text pairs from PubChem. For the single-modal data, we utilize the molecular graph dataset ZINC20 (Irwin et al. 2020) and biomedical text data from PubMed abstracts (White 2020), totaling nearly 100 million entries. Due to computational budget constraints, we are unable to scale up the training size and batch size of the pretraining corpus to the extent achieved by BioT5 (Pei et al. 2023), which utilizes non-open-source pretraining data.
Generation Dataset
For molecule generation tasks, we utilize the ChEBI-20 dataset (Edwards, Zhai, and Ji 2021) for finetuning and evaluation, which consists of 33,010 entries, with 10% allocated for validation and 10% for testing. We adhere to the original data split settings.
Editing Dataset
For molecule editing, we select the retrosynthesis and forward reaction prediction tasks from Mol-Instructions (Fang et al. 2023). The retrosynthesis task begins with a target compound and identifies potential precursor molecules as outputs. The forward reaction prediction task involves forecasting likely products as output molecular graphs, given specific reactants and reagents. Each of these two datasets contains approximately 120,000 reactant-product pairs for training and 1,000 pairs for testing, with each entry accompanied by an instruction.
Figure 3: Some generation examples from our model and baseline models.
model 1-64 65-96 97-128 129-160 $>$ 160
BioT5 (reproduce) 0.810 0.866 0.859 0.817 0.788
@@
Table 3: Analysis for scalability on instruction length.
@@
Several prior studies have explored sequence-based models for evaluation. For the ChEBI-20 dataset, the baseline models include (1) specialist auto-regressive models such as T5 (Raffel et al. 2020), MolT5 (Edwards et al. 2022) and BioT5 (Pei et al. 2023); (2) LLMs such as ChatGPT-3.5 for zero-shot and 10-shot; and (3) text diffusion model tgm-dlm (Gong et al. 2024). For the editing datasets, the baseline models include (1) specialist auto-regressive models such as BioT5 (Pei et al. 2024) and TEXT+CHEM T5 (Christofidellis et al. 2023); (2) LLM such as InstructMol (Cao et al. 2023) and other models. Since BioT5’s pretraining data is not open-sourced, we reproduce their experiments training on our pretraining datasets using reported hyperparameters, rather than referencing their reported performance.
These baselines help us highlight the deficiencies of sequence-based models and clarify our improvements. Additionally, to assess the effectiveness of our diffusion method, we finetuned MoMu-S (Su et al. 2022) on the CHEBI-20 dataset for instruction-based generation. Due to the lack of capability to interpret language in the graph-only structure-aware network, there’re few baselines related to graph diffusion and it’s challenging to directly adapt them into these tasks. Therefore, we don’t consider them as baselines, but further discuss these methods in Ablation Study.
Type Model MACCS RDK Morgan Exact Valid
LLM LLAMA 0.029 0.018 0.017 0.000 0.010
GALACTICA 0.274 0.167 0.134 0.000 0.986
@@
Instruction-Based Molecule Generation
Results
Our results are shown in Table 1. The evaluation metrics are chosen based on previous studies, including a validity metric and five graph-level similarity metrics, with the exception of text-level similarity metrics such as BLEU. Here, we present the results of a single experiment. Detailed descriptions of these metrics, additional results and variance obtained from different seeds, and details for experiments such as hyperparams are available in the supplementary material C,D. Our experimental findings are as follows:
Our model outperforms all auto-regressive methods across almost all similarity metrics with the same level of pretraining corpus. UTGDiff exhibits substantial improvements in the Exact match score, with gains of at least 6% over the best auto-regressive baseline. Also, UTGDiff achieves the highest FTS score, surpassing baseline models by at least 2%. These results indicate the methodological advantages of diffusion paradigm, since it is better suited to capturing the complex structures inherent in molecules.
Compared to text diffusion models, UTGDiff also performs better in both similarity and validity metrics. Especially, UTGDiff exhibits a remarkable 15% improvement in the Exact match score. This indicates that the superior performance of our model is not solely attributable to the diffusion paradigm. Instead, by incorporating graph-based representations, our model leverages explicit structural information in graph data processing, which is not attainable with text diffusion methods.
Type Model MACCS RDK Morgan Exact Valid
LLM LLAMA 0.002 0.001 0.001 0.000 0.010
GALACTICA 0.127 0.036 0.051 0.000 0.995
@@ Results
Our model is far better than the graph-based MoMu-S model. The low performance of MoMu-S may be attributed to the limitations of MoFlow, as it struggles to follow complex instructions and generate long molecules, showing the importance of unified diffusion.
Our model has fewer parameters than other models with comparable performance. Our model has only 125M parameter size, smaller than other T5 models and LLMs.
Our model performs marginally worse than the reported results of BioT5+ and the recently finetuned LLM (Li et al. 2024). This can be attributed to the larger pretraining corpora they used as discussed in the ablation study. Also, our model shows minor shortcomings in validity metrics due to inevitable valence-induced invalidity. The low value in FCD can be attributed to misleading influences from other generated molecules, as discussed in supplementary material A.
Case Study
We present several generated examples in Figure 3. The model successfully generates accurate results comparable to baseline models (Example 1). In some instances, it produces the correct solution when other models fail (Example 2). Also, it can effectively capture essential prompt words like ”tosylate” and generate the required sulfur atom, demonstrating its capability to produce contextually appropriate outputs (Example 3).
Scalability on Instruction
We also present an experiment assessing performance with varying complexities of instructions. We report the MACCS results for varying lengths of instructions and compare them with BioT5. The results indicate that our diffusion model achieves better outcomes across all intervals, demonstrating enhanced scalability when handling more complex instructions.
Instruction-Based Molecule Editing
We also conducted experiments on molecule editing tasks to demonstrate the model’s generalization capabilities across diverse tasks. Transferring from molecular generation to editing requires an additional source molecular graph as input, which is incorporated into the instruction input in graph format. The results are presented in Table 4 and Table 5.
@@
BioT5 - 0.843 1.000 0.763
Table 6: Ablation study on diffusion framework. Steps = 1 is equal to without diffusion
@@
w.o. unifying 0.583 28.47 0.753 -
3M-Diffusion - - - 0.871
UTGDiff 0.867 0.923 0.856 0.957
Table 7: Ablation study on unifying denoising network.
Our model outperforms baseline models in all fingerprint and exact match metrics, with gains of at least 2% in forward reaction prediction tasks. These results underscore the methodological advantages of our model over previous string-based auto-regressive models. When compared to InstructMol-G and InstructMol-GS, our model exhibits a remarkable doubling of the exact match score in forward reaction prediction, with at least a 5% improvement in all metrics across both tasks. These results highlight the significance of the diffusion generation method for these tasks, as InstructMol-G incorporates a 2D molecular graph encoder while retaining a sequence-based decoder approach.
Ablation Study
In our ablation study, we validate our model focusing on the impact of pretraining and model design. All ablation experiments were performed using the CHEBI-20 dataset.
Model Design
Our investigation of model design focuses on two key components to highlight the importance of novel modules: (a) Ablation of the diffusion module. Using the $x_{0}$-parameterization, we can perform inference with flexible step size during the reverse process, while the one-step logits represent a scenario without diffusion. The effectiveness of the diffusion module is demonstrated in Table 6. (b) Ablation of the unified denoising network. This experiment involves the independent use of the graph diffusion network and text encoder, guided by a similarity gradient (Liu et al. 2023b), without the unification of the network. Additionally, we draw from the recent 3M-Diffusion model (Zhu, Xiao, and Honavar 2024), a text-guided graph diffusion method that employs separate encoders, and evaluate our generation results using their Similarity metrics. The results, shown in Table 7, indicate the effectiveness of the unified network.
@@
Pair-only 0.870 0.892 0.879 1.99e-3 300K
All 0.885 0.866 0.893 7.69e-4 100M
@@
Pretraining
We conducted another ablation experiment to assess the impact of pretraining scale. The comparison, detailed in Table 8, highlights the critical roles of both single-modal and paired pretraining datasets and demonstrates the importance of scale, particularly concerning the validity.
Conclusions and Limitations
In this paper, we propose UTGDiff, an instruction-based framework that unifies graph and text data within a single transformer through attention bias, leveraging for the denoising generation of graph data. Our model is evaluated on tasks involving instruction-based molecule generation and editing, showing superior performance while requiring fewer parameters with the same level of pretraining corpus.
Several limitations constrain this research. A critical constraint is our inability to scale the pretraining corpus to a larger size, a factor that significantly benefits models like BioT5 and may have limited our model’s ability to generalize across different molecular structures. Additionally, further experiments on more advanced discrete diffusion frameworks should be explored.
Supplementary Material
A. Analysis of FCD

### v2

@@ 3.4 Reverse Process and Denoising Network
Table 1: Results of the instruction-based molecule generation on ChEBI-20 for both with/without pretraining setting.
Attention Bias
Next, we address the representation of edges. In the attention layer with parameters for values, keys, queries, and outputs, i.e., $W_{V},W_{Q},W_{K}\in\mathbb{R}^{d_{h}\times d_{k}},W_{O}\in\mathbb{R}^{d_{k}\times d_{h}}$, we introduce an additional bias term to incorporate edge information. The attention score between the $i$-th and $j$-th graph tokens is now formalized as:
$\begin{split}\hat{A}_{i,j}^{l}=\frac{1}{\sqrt{d_{k}}}\left(h_{i}^{l}W_{Q}\right)\left(h_{j}^{l}W_{K}\right)^{\top}+b_{i,j}^{l},\\ A^{l}=\mathrm{softmax}(\hat{A}^{l}),\ \text{Attn}(H)=A^{l}H^{l}W_{V}W_{O}\end{split}$ (9)
Here, $b_{i,j}$ denotes the bias added to the attention scores between graph tokens to incorporate structural information [54, 57]. To ensure the expressive power in distinguishing neighbors as one-layer GNNs and the symmetry between edge representations, we therefore represent the initial embedding from the edge category as $b_{i,j}^{0}$ and the update of attention bias as follows:
[…]
The FFN module applied to attention scores follows the structure in standard transformer, including a residual connection and LayerNorm. Our design distinguishes neighbors using edge-specific initial biases $b_{i,j}^{0}$, and enforces symmetry by symmetrizing the attention scores. Beyond structural constraints, our model minimally modifies standard transformer, enabling joint text-graph processing without additional GNNs and allowing full reuse of pre-trained language model weights. Also, layer-wise text-graph interactions enable graph tokens to selectively attend to text tokens and text representations to be refined by graph, supporting fine-grained interactions.
Output
After obtaining the final hidden states $h^{L}$ for tokens and $b^{L}$ for edges through stacked attention layers, we map them back to the vocabulary space for the probability of each category through two separate masked language model heads at the end of the transformer. This yields the node logits $\log p_{\theta}(v_{i}\mid G^{t},S)$ and edge logits $\log p_{\theta}(e_{i,j}\mid G^{t},S)$ for the clean graph distribution. Assembling them into matrices $\hat{p}(V)\in\mathbb{R}^{m\times a}$ and $\hat{p}(E)\in\mathbb{R}^{m\times m\times b}$, the unified denoising network is summarized as:
$\phi_{\theta}(G^{t},S)=\hat{p}(G)=[\hat{p}(V),\hat{p}(E)]$ (11)
Sampling
Empirically, positional embeddings can improve performance in the unified transformer by atom locations, while inherently break permutation equivariance. Therefore, inspired by permutation-based graph encoding methods [37, 39], we adopt a sampling strategy that preserves equivariance at inference: positional embeddings are applied to a fixed node order during training, while random permutations of position indices are introduced during sampling.
Therefore, for any permutation $\pi$, the model generates a graph with node $V$, adjacency matrix $A$, and position indices $i$ with equal probability, satisfying $\mathbb{P}_{i}(V,A)=\mathbb{P}_{\pi^{T}i}(\pi^{T}V,\pi^{T}A\pi)$, thereby ensuring equal probability over the molecule’s permutation set during sampling. We prove the equivariance in Appendix B in the supplementary material [52].
3.5 Training Objective
To train the denoising network, the original method [47] optimizes the cross-entropy loss (denoted as $\mathrm{CE}$) between the predicted probabilities $\hat{p}(G)$ and the clean graph $G$ in the context of single-modal graph diffusion. However, in our multimodal setting, such graph-only supervision is insufficient. Since the model is required to interpret textual instructions while generating molecular structures, an additional loss term related to the text modality is necessary to preserve its language understanding capabilities.
Therefore, we introduce a novel additional textual loss by applying the cross-entropy objective between the predicted logits $\hat{p}(S)$ and the ground-truth instruction tokens. Through the denoising transformer, the logits $\hat{p}(S)$ can obtained alongside $\hat{p}(V)$ and $\hat{p}(E)$ via the masked language model head, formulated as $\phi_{\theta}(G^{t},S)=[\hat{p}(S),\hat{p}(V),\hat{p}(E)]$. Then we define the final training loss $\mathcal{L}$ as:
$\begin{split}\mathcal{L}&=l(\hat{p}(V),V)+l(\hat{p}(E),E)+l(\hat{p}(S),S)\\ \!&=\!\sum_{1\leq i\leq m}\!\mathrm{CE}(v_{i},\!\hat{p}_{i}(V))\!+\!\sum_{1\leq i,j\leq m}\!\mathrm{CE}(e_{i,j},\!\hat{p}_{i,j}(E))\\ &+\!\sum_{1\leq i\leq n}\!\mathrm{CE}(S_{i},\!\hat{p}_{i}(S))\end{split}$ (12)
Model Validity $\uparrow$ Distribution Learning Condition Control
(w/o rule checking) Coverage $\uparrow$ Diversity $\uparrow$ Similarity $\uparrow$ Distance $\downarrow$ Synth. $\downarrow$ O2Perm $\downarrow$ N2Perm $\downarrow$ CO2Perm $\downarrow$ Avg. MAE $\downarrow$
JTVAE-BO 1.0000 (N.A.) 10/11 0.7366 0.7294 23.5990 1.0714 1.0781 1.2352 1.0978 1.1206
DiGress 0.9913 (0.2362) 11/11 0.9099 0.2724 22.7237 2.9842 1.7163 2.0630 1.6738 2.1093
GDSS 0.9205 (0.9076) 9/11 0.7510 0.0000 34.2627 1.3701 1.0271 1.0820 1.0683 1.1369
Graph-DiT-LCC 0.9753 (0.8437) 11/11 0.8875 0.9560 7.0949 1.3099 0.8001 0.9562 0.8125 0.9697
Graph-DiT 0.8245 (0.8437) 11/11 0.8712 0.9600 6.6443 1.2973 0.7440 0.8857 0.7550 0.9205
UTGDiff (Ours) 0.9596 (0.9085) 11/11 0.8532 0.9637 6.0923 1.0208 0.6462 0.7826 0.6459 0.7739
Table 2: Multi-Conditional Generation of 10K Polymers: Results on the synthetic score (Synth.) and three numerical properties (gas permeability for O2, N2, CO2). MAE is calculated between the input conditions and the properties of the generated polymers using Oracles. Best results are highlighted.
This loss helps preserves the model’s language comprehension by reinforcing masked language modeling behavior. When a text token is provided as input, the model is explicitly trained to predict the same token, ensuring token-level consistency. This aligns with the core objective of MLM: reconstructing masked tokens from context demonstrates semantic comprehension, while preserving the identities of unmasked tokens stabilizes the language representations and supports the retention of language understanding during training.
3.6 Pretraining Method
@@
We have mentioned that BERT can be interpreted as a one-step [MASK] absorbing diffusion model which discussed in [2]. This shared training objective enables initialization from pre-trained language models [18]. Accordingly, we initialize our model from the pre-trained RoBERTa [34] to preserve language comprehension.
@@
However, RoBERTa initialization is insufficient for masked prediction on novel graph tokens. To address this, we further pretrain the model on masked graph data while preserving its text modeling ability. We collect paired and single-modal data from textual and graph modalities, randomly mask 15% of tokens and edge indices, and train the model with masked language modeling. The effectiveness of this method on graph data has been validated by prior work [20].
@@
Table 3: Review of pretraining datasets.
Specifically, for multimodal data, we use the PubChem-324K [35], containing 320K molecule-text pairs. For single-modal data, we employ molecular graphs from ZINC20 [22] and biomedical text from PubMed abstracts [50], totaling nearly 100 million entries. Since PubChem-324K is explicitly curated to avoid overlap in generation and others are single-modal data, there’re no risk of data leakage.
4 Experiment
4.1 Instruction-based Molecule generation
First, we evaluate UTGDiff on molecule generation tasks, to demonstrate whether UTGDiff achieves high instruction fidelity and structural validity under structural descriptions.
Dataset
We utilize the ChEBI-20 dataset [11] for fine-tuning and evaluation. It contains 33,010 molecule-instruction pairs, with 10% allocated for validation and 10% for testing.
Baselines
Baseline models include (1) Sequence-based models—T5, MolT5 [12] and BioT5 [41]; (2) Large language models—ChatGPT-3.5, GPT-4 (2024-04-09), and LlaSMol-Mistral-7B fine-tuned on molecular data; (3) text diffusion model tgm-dlm [15]; and (4) graph-based models—MoMu-S [46] (flow-based graph decoder), Digress[47] (predictor-guided diffusion), Graph-DiT [31] (predictor-free multi-condition diffusion), and 3M-Diffusion [59] (separate encoders/decoders). BioT5 is retrained on our pretraining data for fair comparison (Table 3) due to unavailable original pretraining data.
Evaluation
We adopt 6 metrics to assess the similarity and quality of generated molecules: (1) Exact match (Exact); (2) Molecular validity (Validity); (3) 3 fingerprint metrics for similarity in functional group level (FTS); (4) Fréchet ChemNet Distance for similarity in set level (FCD). The full description is in Appendix C in the supplementary material [52].
Results
The results are summarized in Table 1. Additional results with different random seeds and hyperparameter details are provided in Appendix C and D in the supplementary material [52]. Our main findings are as follows:
UTGDiff outperforms all text-based autoregressive methods on similarity metrics with comparable pretraining data size. It achieves substantial improvements in Exact match and FTS scores, surpassing text baselines by 6% on average, highlighting the diffusion paradigm’s strength in capturing complex molecular structures.
UTGDiff significantly outperforms graph-based methods. Previous graph diffusion approaches struggle with complex conditional generation due to information bottleneck and limited scalability. In contrast, UTGDiff enables finer-grained integration of language and graph modalities through unified model and layer-wise aggregration, leading to substantial improvements in instruction-following ability.
Simple text diffusion methods fall short without deep text-graph integration. UTGDiff improves the Exact match score over 15% compared to text diffusion models, indicating that our gains are not solely attributable to the diffusion framework itself but to the deep structural integration of graphs within the language model.
UTGDiff achieves better performance with significantly fewer parameters. Our model contains only 125M parameters, substantially smaller than other models but with better performance.
UTGDiff is marginally outperformed by BioT5+ and a few fine-tuned LLM [29], which can be attributed to their much larger pretraining corpora. Minor shortcomings in molecular validity stem from inevitable valence violations, while lower FCD can due to the metric’s inability to distinguish cross-instruction mismatches, with further analysis in Appendix A in the supplementary material [52].
Figure 3: Some generation examples comparison on CHEBI-20.
Case Study
Several examples are shown in Figure 3. UTGDiff successfully generates accurate molecules as baseline models (Example 1), and can produce correct structures where other models fail (Example 2). Furthermore, it can capture critical instruction details more effective, such as "carboxylic acid group of glycine", demonstrating its ability for contextually appropriate outputs (Example 3).
Length 1-64 65-96 97-128 129-160 >> 160
BioT5 0.810 0.866 0.859 0.817 0.788
@@ Results
BertzCT <100 100-300 300-800 >800
BioT5 0.765 0.863 0.831 0.857
tgl-dlm 0.757 0.859 0.841 0.882
UTGDiff 0.837 0.898 0.881 0.889
Table 4: Analysis for scalability of the model
Scalability
We further assess UTGDiff’s scalability on instruction and molecule complexities. We report MACCS FTS across instruction length and BertzCT intervals, a topological metric reflecting graph size and branching. Table 5 shows that UTGDiff maintains strong performance across all complexity levels, demonstrating robustness on detailed instructions and complex molecules.
MolT5-base 0.733 0.651 0.621 2.20 0.204 0.761
BioT5-base 0.737 0.646 0.595 1.40 0.242 1.000
tgm-dlm 0.741 0.667 0.612 2.06 0.221 0.765
3M-Diffusion 0.495 0.332 0.242 3.47 0.015 1.000
UTGDiff 0.763 0.675 0.623 2.03 0.386 0.847
Table 5: Analysis for generalization in zero-shot experiment
Generalization
To assess the model’s generalization to unseen instructions, we conduct a zero-shot evaluation on PCDes using the model trained only on ChEBI. Sentence-BERT confirms that PCDes instructions differ from ChEBI’s, with 0.929 cosine similarity compared to 0.999. UTGDiff performs competitively across similarity metrics and generalizes well without retraining.
4.2 Multi-Property Conditional Molecule Generation
The instruction dataset above mainly focuses on molecular structures. However, real-world molecular generation scenarios often require inverse design, based on specified property conditions. To better reflect this, we evaluate multi-property conditional generation, formulating numerical properties as part of instructions to assess our model’s ability across more practical conditions. We adapt UTGDiff by expressing property constraints textually and introducing separate MLP encoders to map each condition into an input embedding.
Dataset
We utilize the polymer dataset for materials, which includes three numerical gas permeability properties: O2Perm, CO2Perm, and N2Perm. It contains 553 entries, with 20% for validation and 20% for testing following the data split protocol in [31].
@@ 4.2 Multi-Property Conditional Molecule Generation
We compare against several baselines, including the VAE-based JTVAE model with Bayesian optimization (JTVAE-BO), and diffusion-based methods such as predictor-guided GDSS, DiGress [47], and predictor-free Graph-DiT [31]. We follow the baseline setup in Graph-DiT [31], restricting comparisons to graph-based models, as text-based models are not designed for precise numerical control on property-conditioned generation.
Evaluation
For fair comparison, we adopt the same evaluation protocol on 10,000 generated molecules. In addition to validity and FCD metrics described above, we have: (1) heavy atom type coverage (Coverage); (2) diversity among examples (Diversity); (3) fragment-based similarity with the reference set (Similarity); (4) MAE between the generated and conditioned molecules, including synthetic accessibility score (Synth.) and MAE for numerical conditions (Property).
Results
Our findings are shown in Table 2 and described as follow:
UTGDiff maintains strong molecular distribution modeling beyond prior diffusion methods. UTGDiff achieves the best performance in both similarity and distribution-level metrics. This is notable since it contains only 11 heavy atom types and 5 condition types, much simpler than the diversity seen during pretraining. These results highlight UTGDiff’s generalization from large-scale structure-focused pretraining to property-conditioned generation.
UTGDiff advances conditional graph diffusion and achieves SOTA multi-property control. Across all three properties, UTGDiff consistently reduces MAE, averaging a 19% improvement over the previous SOTA Graph-DiT. This demonstrates that our layer-wise interaction and dedicated numerical encoding enable better multi-property control, validating UTGDiff’s scalability for more diverse property conditions and reflecting real-world generation scenarios.
Figure 4: Some generation examples comparison on O2-N2-CO2 Perm.
Case Study
Several examples are shown in Figure 4. UTGDiff successfully generates near-identical molecules that match the ground truth and outperform baseline models (Example 1). Moreover, it accurately recovers correct atom types (S,Si) under property constraints where other models fail (Example 2 and 3), highlighting UTGDiff’s precise understanding of numerical conditions.
4.3 Instruction-Based Molecule Editing
We further evaluate UTGDiff on molecule editing tasks to assess generalization in diverse scenarios. Unlike generation, editing requires an additional source graph as input, incorporated into the instruction in same graph format as output. Evaluation follows the same metrics as in instruction-based molecule generation.
Dataset
We select the retrosynthesis and forward reaction prediction tasks from Mol-Instructions [13]. The retrosynthesis task predicts precursors from a target compound, while the forward reaction prediction task forecasts products from given reactants and reagents. Each dataset contains 120K training pairs and 1,000 testing pairs.
Baseline
For the editing datasets, the baseline models include (1) specialist auto-regressive models such as BioT5 [41] and TEXT+CHEM T5 [5]; (2) LLM such as InstructMol [4] and others. We also reproduce BioT5’s experiments on our pretraining datasets rather than referencing their reports as discussed above.
Results
The results are presented in Table 6 and Table 7.
[…]
LLM GALACTICA 0.274 0.167 0.134 0.000 0.986
@@ Results
[…]
LLM GALACTICA 0.127 0.036 0.051 0.000 0.995
@@
UTGDiff consistently outperforms all baselines across fingerprint similarity and exact match metrics, achieving 5% improvement in average in two tasks. These results underscore the methodological advantages of our diffusion approach over autoregressive baselines.
Compared to InstructMol-G and InstructMol-GS, we achieves a twofold increase in exact match scores on forward reaction prediction, along with at least a 5% improvement across all metrics on both tasks. This highlights the effectiveness of graph diffusion generation, as InstructMol-G integrates a 2D graph encoder but still relies on a sequence-based decoder, limiting its structural modeling capacity.
4.4 Ablation Study
We ablate our model on the model design and training strategy on CHEBI-20 dataset, validating the contribution of these key designs.
Model Design
We conduct ablation studies focusing on two aspects of model design: (1) the contribution of the diffusion module to generation performance; and (2) the effectiveness of key architectural components, including unification via attention bias and the absorbing state implemented through [MASK] tokens.
@@ 4.4 Ablation Study
Table 8: Ablation on diffusion. Steps = 1 means without diffusion.
(1) Effect of the Diffusion Module. The $x_{0}$-parameterization allow us to use flexible step-size inference during generation. Performing generation with one-step logits represents a setting without diffusion. As shown in Table 8, the diffusion module improves performance, demonstrating the importance of progressive refinement.
@@
Cross-attn 0.693 2.720 0.632
marginal absorbing 0.877 1.852 0.837
UTGDiff 0.885 0.866 0.893
Table 9: Ablation on our model design.
(2) Effect of Architecture. To ablate the unification strategy, we replace attention bias with cross-attention, encoding instructions separately for cross-modal fusion. To ablate the absorbing state design, we replace the [MASK]-based noise representation with marginal distributions as Graph-DiT [31]. Table 9 confirms the effectiveness of both attention bias and [MASK]-based noise modeling.
@@
no textual loss 0.882 1.224 0.833 3.57e-3 100M
no position index 0.764 2.387 0.821 2.92e-1 100M
@@
Training Strategy
We conducted ablation experiments to assess the impact of pretraining scale and the textual loss $l(\hat{p}(S),S)$ during fine-tuning. Table 10 highlights the importance of both single-modal and paired pretraining datasets, demonstrating that scaling pretraining data significantly benefits downstream performance. The comparison of textual loss shows that incorporating textual loss leads to better convergence and higher validity. The comparison of position index shows the necessity of our sampling stratage.
5 Conclusions and Limitations
We propose UTGDiff, a unified instruction-based framework that integrates graph and text within a single transformer. With absorbing noise, atom-specific tokens, and attention bias, it enables fine-grained graph diffusion generation conditioned on complex instructions. Our model excels in molecule generation and editing, covering instructions for molecular structures or properties, while maintaining a smaller parameter size compared to existing baselines.
Despite these advancements, UTGDiff is limited by the scale of its pretraining corpus, which may constrain generalization to more diverse molecular structures. Future work will focus on scaling pretraining and exploring more advanced discrete diffusion techniques to further enhance instruction-conditioned generation.
Appendix A Analysis of FCD